OpenStudy (anonymous):

MEDAL FOR RIGHT ANSWER!!! PLEASE HELP!!! What is $144.99 after it is discounted by 15%? SHOW YOUR WORK

OpenStudy (anonymous):

@horsegirl27 @lexie<3

OpenStudy (horsegirl27):

i can help:)

OpenStudy (anwaarullah):

144.99 * .15= 21.75 144.99-21.75= 123.24

OpenStudy (alexandervonhumboldt2):

144.99*(1-15/100)

OpenStudy (alexandervonhumboldt2):

123.2415

OpenStudy (horsegirl27):

So first we need to find the discount

OpenStudy (horsegirl27):

do that^^^

OpenStudy (anonymous):

I need to know the formula to find it

OpenStudy (horsegirl27):

You need to find the discount, then subtract it from $144.99 to get the discounted price

OpenStudy (horsegirl27):

ok I can help with that

OpenStudy (alexandervonhumboldt2):

i told you it 144.99*(1-15/100) is the final formula

OpenStudy (anonymous):

I don't get it though, 1-15?

OpenStudy (anonymous):

123.24 all you have to do is put it in the calculator

OpenStudy (alexandervonhumboldt2):

1-15/100=0.85

OpenStudy (alexandervonhumboldt2):

0.85*144.99=123.2415

OpenStudy (anwaarullah):

Discount= List price* discount rate